# Variational formulations of transport on combinatorial meshes

### v6

@@ 2.2 Metric operations
In our applications $M$ is a space manifold and so it is useful to assume that $g$ has physical dimensions $\mathsf{L}^{2}$. Indeed, let $x\in M$ and $v\in T_{x}M$ be a tangent vector, $\left.g\right|_{x}$ be the fibre of $g$ at $x$. Then the length of $v$ is defined as $\sqrt{\left.g\right|_{x}(v,v)}$, and it is natural for it to have physical dimension $[[v]]\cdot\mathsf{L}$. Hence, $g_{x}$, and $g$ as well, is of physical dimension $\mathsf{L}^{2}$.
@@
$[[g^{*}_{D-p}]]\,[[\star_{p}]]\,[[\mathop{\rm vol}\nolimits]]=1\Rightarrow\mathsf{L}^{-2(D-p)}\,[[\star_{p}]]\,\mathsf{L}^{D}=1\Rightarrow[[\star_{p}]]=\mathsf{L}^{D-2p}.$ (2.58)
@@
Consider an oriented orthonormal basis $(e^{1},...,e^{D})$ of $\Omega^{1}M$, $\mathop{\rm vol}\nolimits:=e^{1}\wedge...\wedge e^{D}$ be the volume form. Let $I=(I_{1},...,I_{p})$, $1\leq I_{1}<...<I_{p}\leq D$ and $e_{I}:=e^{I_{1}}\wedge...\wedge e^{I_{p}}$. Let $J$ be the ordered complement of $I$ (with respect to $(1,...,D)$), and $e^{J}:=e^{J_{1}}\wedge...\wedge e_{J_{D-p}}$. Then $\star_{p}e^{I}=se^{J}$ for some $s\in\{-1,1\}$ (since taking the wedge product gives $e^{1}\wedge...\wedge e^{D}=\mathop{\rm vol}\nolimits$ after reordering, taken into account by $s$), and similarly $\star_{D-p}e^{J}=te^{I}$ for some $t\in\{-1,1\}$. Then
@@
$[[\langle\cdot,\cdot\rangle_{p}]]=[[d^{\star}_{p}]]\,[[\langle\cdot,\cdot\rangle_{p-1}]]\Rightarrow\mathsf{L}^{D-2p}=[[d^{\star}_{p}]]\,\mathsf{L}^{D-2(p-1)}\Rightarrow[[d^{\star}_{p}]]=\mathsf{L}^{-2}.$ (2.72)
@@ 3 Exterior calculus variational formulations
is characterised as follows: for any $D$-dimensional subregion $V$ of $M$, and any time interval $[t_{1},t_{2}]\subset I$,
[…]
@@
is characterised as follows. Let $[t_{1},t_{2}]\subset I$ be a time interval, $S$ be a $(D-1)$-dimensional hypersurface on $M$. Consider a subregion $V$ of $M$ partitioned into two parts $V_{+}$ and $V_{-}$ sharing $S$ as a common boundary, with relative orientations $\varepsilon(V_{+},S)=1$ and $\varepsilon(V_{-},S)=-1$. Define the net flow of the amount $A$ through $S$ in $[t_{1},t_{2}]$ to be be the difference between the flow from $V_{+}$ to $V_{-}$ in $[t_{1},t_{2}]$ and the flow from $V_{-}$ to $V_{+}$ in $[t_{1},t_{2}]$. Then
[…]
@@
In Equation 3.6 the perspective is from the point of view of the hypersurface $S$. However, if we view it through the eyes of $V_{+}$, the aforementioned term is the net outflow of $A$ from $V_{+}$ to $V_{-}$ through $S$ in $[t_{1},t_{2}]$, while from the perspective of $V_{-}$ it is the net inflow of $A$ in $V_{-}$ from $V_{+}$ through $S$ in $[t_{1},t_{2}]$.
@@
Let $V$ be a subregion of $M$, $[t_{1},t_{2}]\subset I$. Then
@@
[…]
@@
is characterised as follows: for any $D$-dimensional subregion $V$ of $M$, and any time interval $[t_{1},t_{2}]\subset I$,
[…]
@@
The conservation law for an extensive quantity $A$ (e.g., energy) has the following intuitive formulation: for any $D$-dimensional subregion $V$ of $M$, and any time interval $[t_{1},t_{2}]\subset I$,
@@
[…]
Since $V$ and $[t_{1},t_{2}]$ are arbitrary, we can remove the integrals over them. Thus, we arrive at the differential formulation of the conservation law:
@@
Potentials for mass (mass transfer), energy (heat transfer), charge (charge transport), volume (volume transport) are concentration, temperature, electric potential, and pressure respectively.
@@
In heat transfer the volumetric capacity is the volumetric heat capacity; in charge transport the volumetric capacity is the volumetric capacitance; in mass and volume transport the potentials are already the densities of the extensive quantities and so the respective volumetric capacities are not presented (they are identically $1$).
@@
The conductivities for mass, heat energy, charge and volume transport are mass diffusivity, thermal conductivity, electrical conductivity and hydraulic conductivity. The respective constitutive laws are Fick’s law of diffusion, Fourier’s law, Ohm’s law, and Darcy’s law.
@@
Phenomenon Amount $[\mathsf{X}]$ Potential $[\mathsf{Y}]$ Volumetric capacity $[\mathsf{X}\mathsf{Y}^{-1}\mathsf{L}^{-D}]$ Conductivity $[\mathsf{X}\mathsf{Y}^{-1}\mathsf{L}^{2-D}\mathsf{T}^{-1}]$ Constitutive law
Mass transport Mass $[\mathsf{M}]$ Mass density $[\mathsf{M}\mathsf{L}^{-D}]$ 1 $[1]$ Mass diffusivity $[\mathsf{L}^{2}\mathsf{T}^{-1}]$ Fick’s law
Heat transport Heat energy $[\mathsf{M}\mathsf{L}^{2}\mathsf{T}^{-2}]$ Temperature $[\theta]$ Volumetric heat capacity $[\mathsf{M}\mathsf{L}^{2-D}\mathsf{T}^{-2}\theta^{-1}]$ Thermal conductivity $[\mathsf{M}\mathsf{L}^{4-D}\mathsf{T}^{-3}\theta^{-1}]$ Fourier’s law
Charge transport Charge $[\mathsf{C}]$ Electric potential $[\mathsf{M}\mathsf{L}^{2}\mathsf{T}^{-2}\mathsf{C}^{-1}]$ Volumetric capacitance $[\mathsf{M}^{-1}\mathsf{L}^{-2-D}\mathsf{T}^{2}\mathsf{C}^{2}]$ Electrical conductivity $[\mathsf{M}^{-1}\mathsf{L}^{-D}\mathsf{T}\mathsf{C}^{2}]$ Ohm’s law
Fluid flow through porous media Volume $[\mathsf{L}^{D}]$ Pressure $[\mathsf{M}\mathsf{L}^{2-D}\mathsf{T}^{-2}]$ Compressibility $[\mathsf{M}^{-1}\mathsf{L}^{D-2}\mathsf{T}^{2}]$ Hydraulic conductivity $[\mathsf{M}^{-1}\mathsf{L}^{D}\mathsf{T}]$ Darcy’s law
@@
be the volumetric flow rate pseudo-form, which is a prescribed quantity, e.g., calculated from a volume transport problem. It is characterised by the following property: for any two subregions $V_{+}$ and $V_{-}$ with common boundary $S$ (and respective positive and negative relative orientations with $S$), and any time interval $[t_{1},t_{2}]\subset I$,
@@
Amount $Q$ $\mathcal{C}^{\infty}(I,\Omega^{D}M)$ $\mathsf{X}$ Yes eq. 3.3
Flow rate $q$ $\mathcal{C}^{\infty}(I,\Omega^{D-1}M)$ $\mathsf{X}\mathsf{T}^{-1}$ Yes eq. 3.5
Diffusive flow rate $q_{\mathcal{D}}$ $\mathcal{C}^{\infty}(I,\Omega^{D-1}M)$ $\mathsf{X}\mathsf{T}^{-1}$ Yes eq. 3.24
Advective flow rate $q_{\mathcal{A}}$ $\mathcal{C}^{\infty}(I,\Omega^{D-1}M)$ $\mathsf{X}\mathsf{T}^{-1}$ Yes eq. 3.25
Potential $u$ $\mathcal{C}^{\infty}(I,\Omega^{0}M)$ $\mathsf{Y}$ No eq. 3.16
@@
$\frac{\partial Q}{\partial t}$ $=f-dq$ $\mathsf{X}\mathsf{T}^{-1}$ Conservation law eq. 3.13
$\frac{\partial Q}{\partial t}$ $=\star_{0}\widetilde{\pi}\frac{\partial u}{\partial t}=\pi\star_{0}\frac{\partial u}{\partial t}$ $\mathsf{X}\mathsf{T}^{-1}$ Relation between potential and amount eq. 3.23
$q$ $=q_{\mathcal{D}}+q_{\mathcal{A}}$ $\mathsf{X}\mathsf{T}^{-1}$ Flow rate decomposition eq. 3.26
$q_{\mathcal{D}}$ $=\kappa d^{\star}_{D}\star_{0}u$ $\mathsf{X}\mathsf{T}^{-1}$ Constitutive law eq. 3.29
$q_{\mathcal{A}}$ $=(\star_{D}Q)\wedge v$ $\mathsf{X}\mathsf{T}^{-1}$ Advective flow rate formula eq. 3.35
$\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{D},0}}(u)$ $=g_{\mathbf{D}}$ $\mathsf{Y}$ Dirichlet boundary condition eq. 3.19
$\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{N},D-1}}(q)$ $=g_{\mathbf{N}}$ $\mathsf{X}\mathsf{T}^{-1}$ Neumann boundary condition eq. 3.15
$u(t_{0})$ $=u_{0}$ $\mathsf{Y}$ Initial condition eq. 3.37
@@ 3.2 Mixed weak formulations of transport phenomena
$A\colon\Omega^{D-1}M\times\mathcal{C}^{\infty}(I,\Omega^{D-1}M)\to\mathbb{R},$ $A(r,s):=\langle r,\kappa^{-1}s\rangle_{M,D-1}$ $[\mathsf{X}^{-1}\mathsf{T}\mathsf{Y}],$ (3.51a)
$B_{\mathcal{D}}\colon\Omega^{D}M\times\mathcal{C}^{\infty}(I,\Omega^{D-1}M)\to\mathbb{R},$ $B_{\mathcal{D}}(\widetilde{w},r):=\langle d_{D-1}r,\widetilde{w}\rangle_{M,D}$ $[\mathsf{L}^{-D}],$ (3.51b)
$B_{\mathcal{A}}\colon\Omega^{D}M\times\mathcal{C}^{\infty}(I,\Omega^{D-1}M)\to\mathbb{R},$ $B_{\mathcal{A}}(\widetilde{w},r):=\langle r,\kappa^{-1}(\star_{D}\pi\widetilde{w}\wedge v)\rangle_{M,D-1}$ $[\mathsf{L}^{-D}],$ (3.51c)
$C\colon\Omega^{D}M\times\mathcal{C}^{\infty}(I,\Omega^{D}M)\to\mathbb{R},$ $C(\widetilde{w},\widetilde{u}):=\langle\pi\widetilde{u},\widetilde{w}\rangle_{M,D}$ $[\mathsf{X}\mathsf{L}^{-2D}\mathsf{Y}^{-1}],$ (3.51d)
$G\colon\Omega^{D-1}M\to\mathbb{R},$ $G(r):=\int_{\Gamma_{\mathbf{D}}}(\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{D}},D-1}r\wedge g_{\mathbf{D}})$ $[\mathsf{Y}],$ (3.51e)
$F\colon\Omega^{D}M\to\mathbb{R},$ $F(\widetilde{w}):=\langle f,\widetilde{w}\rangle_{M,D}$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{L}^{-D}],$ (3.51f)
$\mathtt{flow\_rate}\colon\Omega^{0}M\to\mathbb{R},$ $\mathtt{flow\_rate}(w):=\kappa\,d^{\star}_{D}\star_{0}w+\pi w\wedge v$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{Y}^{-1}].$ (3.51g)
@@
$\forall r[\mathsf{X}\mathsf{T}^{-1}]\in\mathop{\rm Ker}\nolimits\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{N}},D-1},\;$ $A(r,q)-B_{\mathcal{D}}^{T}(r,\widetilde{u})-B_{\mathcal{A}}^{T}(r,\widetilde{u})$ $=-G(r)\;$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{Y}],$ (3.52a)
$\forall\widetilde{w}[\mathsf{Y}\mathsf{L}^{D}]\in\Omega^{D}M,\;$ $-B_{\mathcal{D}}(\widetilde{w},q)-C(\widetilde{w},\frac{\partial\widetilde{u}}{\partial t})$ $=-F(\widetilde{w})\;$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{Y}],$ (3.52b)
@@
$\forall r[\mathsf{X}\mathsf{T}^{-1}]\in\mathop{\rm Ker}\nolimits\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{N}},D-1},\;$ $A(r,q)-B_{\mathcal{D}}^{T}(r,\widetilde{u})-B_{\mathcal{A}}^{T}(r,\widetilde{u})$ $=-G(r)\;$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{Y}],$ (3.54a)
$\forall\widetilde{w}[\mathsf{Y}\mathsf{L}^{D}]\in\Omega^{D}M,\;$ $-B_{\mathcal{D}}(\widetilde{w},q)$ $=-F(\widetilde{w})\;$ $[\mathsf{X}\mathsf{T}^{-1}\mathsf{Y}],$ (3.54b)

### v7

@@ 2.2 Metric operations
In our applications $M$ is a space manifold and so it is useful to assume that $g$ has physical dimensions $\mathsf{L}^{2}$. Indeed, let $x\in M$ and $v\in T_{x}M$ be a tangent vector, $\left.g\right|_{x}$ be the fibre of $g$ at $x$. Then the length of $v$ is defined as $\sqrt{\left.g\right|_{x}(v,v)}$, and it is natural for it to have physical dimension $v]]\cdot\mathsf{L}$. Hence, $g_{x}$, and $g$ as well, is of physical dimension $\mathsf{L}^{2}$.
@@
$g^{*}_{D-p}]]\,[[\star_{p}]]\,[[\mathop{\rm vol}\nolimits]]=1\Rightarrow\mathsf{L}^{-2(D-p)}\,[[\star_{p}]]\,\mathsf{L}^{D}=1\Rightarrow[[\star_{p}]]=\mathsf{L}^{D-2p}.$ (2.58)
@@
Consider an oriented orthonormal basis $(e^{1},...,e^{D})$ of $\Omega^{1}M$, $\mathop{\rm vol}\nolimits:=e^{1}\wedge...\wedge e^{D}$ be the volume form. Let $I=(I_{1},...,I_{p})$, $1\leq I_{1}<...<I_{p}\leq D$ and $e_{I}:=e^{I_{1}}\wedge...\wedge e^{I_{p}}$. Let $J$ be the ordered complement of $I$ (with respect to (1,,D)(1,...,D)), and $e^{J}:=e^{J_{1}}\wedge...\wedge e_{J_{D-p}}$. Then $\star_{p}e^{I}=se^{J}$ for some $s\in\{-1,1\}$ (since taking the wedge product gives $e^{1}\wedge...\wedge e^{D}=\mathop{\rm vol}\nolimits$ after reordering, taken into account by $s$), and similarly $\star_{D-p}e^{J}=te^{I}$ for some $t\in\{-1,1\}$. Then
@@
$\langle\cdot,\cdot\rangle_{p}]]=[[d^{\star}_{p}]]\,[[\langle\cdot,\cdot\rangle_{p-1}]]\Rightarrow\mathsf{L}^{D-2p}=[[d^{\star}_{p}]]\,\mathsf{L}^{D-2(p-1)}\Rightarrow[[d^{\star}_{p}]]=\mathsf{L}^{-2}.$ (2.72)
@@ 3 Exterior calculus variational formulations
is characterised as follows: for any $D$-dimensional subregion $V$ of $M$, and any time interval $t_{1},t_{2}]\subset I$,
$\textrm{``amount difference of $A$ inside $V$ in $[t_{1},t_{2}]$''}=\int_{V}Q(t_{2})-\int_{V}Q(t_{1}).$ (3.4)
@@
is characterised as follows. Let $t_{1},t_{2}]\subset I$ be a time interval, $S$ be a (D1)(D-1)-dimensional hypersurface on $M$. Consider a subregion $V$ of $M$ partitioned into two parts $V_{+}$ and $V_{-}$ sharing $S$ as a common boundary, with relative orientations $\varepsilon(V_{+},S)=1$ and $\varepsilon(V_{-},S)=-1$. Define the net flow of the amount $A$ through $S$ in $t_{1},t_{2}]$ to be be the difference between the flow from $V_{+}$ to $V_{-}$ in $t_{1},t_{2}]$ and the flow from $V_{-}$ to $V_{+}$ in $t_{1},t_{2}]$. Then
$\textrm{``net flow of $A$ through $S$ in $[t_{1},t_{2}]$''}=\int_{t_{1}}^{t_{2}}\int_{S}q\,dt.$ (3.6)
@@
In Equation 3.6 the perspective is from the point of view of the hypersurface $S$. However, if we view it through the eyes of $V_{+}$, the aforementioned term is the net outflow of $A$ from $V_{+}$ to $V_{-}$ through $S$ in $t_{1},t_{2}]$, while from the perspective of $V_{-}$ it is the net inflow of $A$ in $V_{-}$ from $V_{+}$ through $S$ in $t_{1},t_{2}]$.
@@
Let $V$ be a subregion of $M$, $t_{1},t_{2}]\subset I$. Then
@@
By definition, the boundary $\partial V$ has the induced boundary orientation, i.e., $\varepsilon(V,\partial V)=1,\ \varepsilon(M\setminus\overline{V},\partial V)=-1$. The net outflow of $A$ from $V$ through its boundary in $t_{1},t_{2}]$, hence, equals to the net flow of $A$ through $\partial V$ in $t_{1},t_{2}]$, which by Definition 3.3 is $\int_{t_{1}}^{t_{2}}\int_{\partial V}q\,dt$. ∎
@@
is characterised as follows: for any $D$-dimensional subregion $V$ of $M$, and any time interval $t_{1},t_{2}]\subset I$,
$\textrm{``internal production of $A$ inside $V$ in $[t_{1},t_{2}]$''}=\int_{t_{1}}^{t_{2}}\int_{V}f\,dt.$ (3.9)
@@
The conservation law for an extensive quantity $A$ (e.g., energy) has the following intuitive formulation: for any $D$-dimensional subregion $V$ of $M$, and any time interval $t_{1},t_{2}]\subset I$,
@@
$\int_{t_{1}}^{t_{2}}\left(\int_{V}\frac{\partial Q}{\partial t}\right)\,dt=\int_{t_{1}}^{t_{2}}\left(\int_{V}f\right)\,dt-\int_{t_{1}}^{t_{2}}\left(\int_{V}\,d_{D-1}q\right)\,dt.$ (3.12)
Since $V$ and $t_{1},t_{2}]$ are arbitrary, we can remove the integrals over them. Thus, we arrive at the differential formulation of the conservation law:
@@
Potentials for mass (mass transfer), energy (heat transfer), charge (charge transport), volume (fluid flow through porous media) are concentration, temperature, electric potential, and pressure respectively.
@@
In mass transfer the potential is the mass density, so no volumetric capacity is present (it is given by the number $1$). In heat transfer, charge transport and fluid flow through porous media the volumetric capacities are volumetric heat capacity, volumetric capacitance, and compressibility respectively.
@@
The conductivities for mass transfer, heat transfer, charge transport, and fluid flow through porous media are mass diffusivity, thermal conductivity, electrical conductivity, and hydraulic conductivity respectively. The respective constitutive laws are Fick’s law of diffusion, Fourier’s law, Ohm’s law, and Darcy’s law.
@@
Phenomenon
Amount
$\mathsf{X}]$
Potential
$\mathsf{Y}]$
Volumetric capacity
$\mathsf{X}\mathsf{Y}^{-1}\mathsf{L}^{-D}]$
Conductivity
$\mathsf{X}\mathsf{Y}^{-1}\mathsf{L}^{2-D}\mathsf{T}^{-1}]$
Constitutive
law
Mass transfer
Mass
$\mathsf{M}]$
Mass density
$\mathsf{M}\mathsf{L}^{-D}]$
1
$[1]$
Mass diffusivity
$\mathsf{L}^{2}\mathsf{T}^{-1}]$
Fick’s law
Heat transfer
Heat energy
$\mathsf{M}\mathsf{L}^{2}\mathsf{T}^{-2}]$
Temperature
$\theta]$
Volumetric heat capacity
$\mathsf{M}\mathsf{L}^{2-D}\mathsf{T}^{-2}\theta^{-1}]$
Thermal conductivity
$\mathsf{M}\mathsf{L}^{4-D}\mathsf{T}^{-3}\theta^{-1}]$
Fourier’s law
Charge transport
Charge
$\mathsf{C}]$
Electric potential
$\mathsf{M}\mathsf{L}^{2}\mathsf{T}^{-2}\mathsf{C}^{-1}]$
Volumetric capacitance
$\mathsf{M}^{-1}\mathsf{L}^{-2-D}\mathsf{T}^{2}\mathsf{C}^{2}]$
Electrical conductivity
$\mathsf{M}^{-1}\mathsf{L}^{-D}\mathsf{T}\mathsf{C}^{2}]$
Ohm’s law
Fluid flow through
porous media
Volume
$\mathsf{L}^{D}]$
Pressure
$\mathsf{M}\mathsf{L}^{2-D}\mathsf{T}^{-2}]$
Compressibility
$\mathsf{M}^{-1}\mathsf{L}^{D-2}\mathsf{T}^{2}]$
Hydraulic conductivity
$\mathsf{M}^{-1}\mathsf{L}^{D}\mathsf{T}]$
Darcy’s law
@@
be the volumetric flow rate pseudo-form, which is a prescribed quantity, e.g., calculated from a volume transport problem. It is characterised by the following property: for any two subregions $V_{+}$ and $V_{-}$ with common boundary $S$ (and respective positive and negative relative orientations with $S$), and any time interval $t_{1},t_{2}]\subset I$,
@@
Amount $Q$ $\mathcal{C}^{\infty}(I,\Omega^{D}M)$ $\mathsf{X}$ Yes Equation 3.3
Flow rate $q$ $\mathcal{C}^{\infty}(I,\Omega^{D-1}M)$ $\mathsf{X}\mathsf{T}^{-1}$ Yes Equation 3.5
Diffusive flow rate $q_{\mathcal{D}}$ $\mathcal{C}^{\infty}(I,\Omega^{D-1}M)$ $\mathsf{X}\mathsf{T}^{-1}$ Yes Equation 3.24
Advective flow rate $q_{\mathcal{A}}$ $\mathcal{C}^{\infty}(I,\Omega^{D-1}M)$ $\mathsf{X}\mathsf{T}^{-1}$ Yes Equation 3.25
Potential $u$ $\mathcal{C}^{\infty}(I,\Omega^{0}M)$ $\mathsf{Y}$ No Equation 3.16
@@
$\frac{\partial Q}{\partial t}$ =fdq=f-dq $\mathsf{X}\mathsf{T}^{-1}$ Conservation law Equation 3.13
$\frac{\partial Q}{\partial t}$ $\star_{0}\widetilde{\pi}\frac{\partial u}{\partial t}=\pi\star_{0}\frac{\partial u}{\partial t}$ $\mathsf{X}\mathsf{T}^{-1}$ Relation between potential and amount Equation 3.23
$q$ $q_{\mathcal{D}}+q_{\mathcal{A}}$ $\mathsf{X}\mathsf{T}^{-1}$ Flow rate decomposition Equation 3.26
$q_{\mathcal{D}}$ $\kappa d^{\star}_{D}\star_{0}u$ $\mathsf{X}\mathsf{T}^{-1}$ Constitutive law Equation 3.29
$q_{\mathcal{A}}$ $(\star_{D}Q)\wedge v$ $\mathsf{X}\mathsf{T}^{-1}$ Advective flow rate formula Equation 3.35
$\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{D},0}}(u)$ $g_{\mathbf{D}}$ $\mathsf{Y}$ Dirichlet boundary condition Equation 3.19
$\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{N},D-1}}(q)$ $g_{\mathbf{N}}$ $\mathsf{X}\mathsf{T}^{-1}$ Neumann boundary condition Equation 3.15
$u(t_{0})$ $u_{0}$ $\mathsf{Y}$ Initial condition Equation 3.37
@@ 3.2 Mixed weak formulations of transport phenomena
$A\colon\Omega^{D-1}M\times\mathcal{C}^{\infty}(I,\Omega^{D-1}M)\to\mathbb{R},\$$A(r,s):=\langle r,\kappa^{-1}s\rangle_{M,D-1}\$$[\mathsf{X}^{-1}\mathsf{T}\mathsf{Y}],$ (3.51a)
$B_{\mathcal{D}}\colon\Omega^{D}M\times\mathcal{C}^{\infty}(I,\Omega^{D-1}M)\to\mathbb{R},\$$B_{\mathcal{D}}(\widetilde{w},r):=\langle d_{D-1}r,\widetilde{w}\rangle_{M,D}\$$[\mathsf{L}^{-D}],$ (3.51b)
$B_{\mathcal{A}}\colon\Omega^{D}M\times\mathcal{C}^{\infty}(I,\Omega^{D-1}M)\to\mathbb{R},\$$B_{\mathcal{A}}(\widetilde{w},r):=\langle r,\kappa^{-1}(\star_{D}\pi\widetilde{w}\wedge v)\rangle_{M,D-1}\$$[\mathsf{L}^{-D}],$ (3.51c)
$C\colon\Omega^{D}M\times\mathcal{C}^{\infty}(I,\Omega^{D}M)\to\mathbb{R},\$$C(\widetilde{w},\widetilde{u}):=\langle\pi\widetilde{u},\widetilde{w}\rangle_{M,D}\$$[\mathsf{X}\mathsf{L}^{-2D}\mathsf{Y}^{-1}],$ (3.51d)
$G\colon\Omega^{D-1}M\to\mathbb{R},\$$G(r):=\int_{\Gamma_{\mathbf{D}}}(\mathop{\rm tr}\nolimits_{\Gamma_{\mathbf{D}},D-1}r\wedge g_{\mathbf{D}})\$$[\mathsf{Y}],$ (3.51e)
$F\colon\Omega^{D}M\to\mathbb{R},\$$F(\widetilde{w}):=\langle f,\widetilde{w}\rangle_{M,D}\$$[\mathsf{X}\mathsf{T}^{-1}\mathsf{L}^{-D}],$ (3.51f)
$\mathtt{flow\_rate}\colon\Omega^{0}M\to\mathbb{R},\$$\mathtt{flow\_rate}(w):=\kappa\,d^{\star}_{D}\star_{0}w+\pi w\wedge v\$$[\mathsf{X}\mathsf{T}^{-1}\mathsf{Y}^{-1}].$ (3.51g)
@@
[…]
@@
[…]